# ACEGEN: Reinforcement learning of generative chemical agents for drug discovery

### v1

@@ 3.1 Benchmarking RL performance
To assess different RL algorithms, we have compared performance on the Practical Molecular Optimization (MolOpt) benchmark Gao et al. (2022) as implemented in MolScore (Thomas et al., 2023). This benchmark encompasses 23 distinct tasks associated with different objectives to be optimized within a budget of 10,000 molecules. All algorithms use the same GRU policy model architecture as in the original MolOpt benchmark, implemented in ACEGEN and pretrained on a curated subset of ChEMBL (Gaulton et al., 2012). We compared REINFORCE (Williams, 1992), REINVENT with the hyperparameters from the original paper (Blaschke et al., 2020), REINVENT-MolOpt with optimized hyperparameters for the MolOpt benchmark (Gao et al., 2022), and AHC with the hyperparameters from the original paper (Thomas et al., 2022). Note that we utilize experience replay in all of the aforementioned algorithms Then we also test A2C (Mnih et al., 2016), PPO (Schulman et al., 2017) which does not utilize any experience replay, and PPOD (Libardi et al., 2021) which is an adaptation of PPO to accommodate experience replay. All algorithms are ACEGEN implementations, and the hyperparameter values used for all of them are provided in our repository. As a sanity check, we compared our implementation of REINVENT with the implementation used in the MolOpt paper. Our simplified re-implementation achieved better results with faster training. These results are shown in Appendix C.
As proposed by the MolOpt benchmark authors, Table 1 shows algorithm performance on the AUC of the top 10 molecules as an indication of the sample efficiency of identifying 10 desirable molecules with respect to the objective, a representative figure that might be carried forward to later stages of drug design. This shows that PPOD is state-of-the-art with respect to sample efficiency, followed by REINVENT-MolOpt and then PPO. To achieve a higher level understanding of performance, Table 2 shows a selection of metrics aimed to measure maximum performance, efficiency, and exploration (a common trade-off with exploitation and hence efficiency in RL). These results show that REINVENT-MolOpt achieves maximum performance as measured by the average top 10 molecules, PPOD maximum efficiency as measured by AUC of the top 10 molecules, and AHC maximum exploration as measured by the number of unique compounds generated.
@@
Table 2: Algorithm comparison on a selection of metrics analyzing the performance on the MolOpt benchmark. Valid is a sanity check that measures the proportion of valid molecules generated. Top-10 AUC is a measure of sample efficiency as shown in 1. Top-10 Avg is a measure of the absolute best 10 molecules achieved with the budget. Unique is a proxy for exploration which measures the proportion of unique molecules generated. Basic and target chemistry filters (B&T-CF) are the proportion of molecules after filtering out highly idiosyncratic molecules in general or with respect to the pretraining dataset. The B&T-CF Top-10 AUC and B&T-CF Top-10 Avg are recalculated on this subset while also enforcing that the 10 molecules identified offer diverse solutions (Div). Lastly, B&T-CF Diversity is a proxy measure of exploration via sphere exclusion diversity Thomas et al. (2021) on this subset. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. A perfect score for all metrics is 23.
@@ 3.2 Benchmarking RL performance for practical drug discovery
RL optimizes the policy for the cumulative future rewards as provided by either one or a sum of multiple scoring functions. However, in practical drug discovery, it is not straightforward to write the precise scoring functions that are needed and how to weigh them into a single scalar. For example, scoring functions can sometimes not produce the desired chemistry, or show clear exploitation loopholes Renz et al. (2019); Langevin et al. (2022). In this setting, a sufficient element of regularization to a prior policy is useful, as the prior policy has learned a chemical space distribution based on a specified training dataset of choice which acts as a representation of desirable chemistry by examples. It is for this reason that methods like REINVENT have been designed heuristically to contain terms that try to enforce regularization instead of entirely using the reward signal.
To account for this, we make a modification to the average and AUC of the top 10 molecules to ensure they are diverse by an ECFP4 Tanimoto similarity of less than 0.35 to each other. Moreover, we use sphere exclusion diversity (SEDiv@1k) of de novo molecules Thomas et al. (2021) which measures the proportion of molecules needed to describe chemical space in a random sample of 1,000 molecules, as a proxy for exploration. This is more representative than the number of unique molecules that could all reside in a close area of chemical space. Lastly, we remove de novo molecules that do not pass a series of filters. This includes a basic chemistry filter (B-CF): logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in Polykovskiy et al. (2020). Moreover on the assumption that the training dataset describes desirable chemistry by examples, a target chemistry filter (T-CF): logP and molecular weight within $\mu\pm 4\sigma$ of the training dataset distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). The combination of the basic and target chemistry filter is denoted (B&T-CF).
@@ 3.3 Ablation study of the REINVENT algorithm
One of the most common RL algorithms used for the fine-tuning of CLMs is REINVENT Olivecrona et al. (2017). From a theoretical RL perspective, this algorithm can be viewed as a reward shaping of the REINFORCE algorithm. However, the derivation of this specific reward-shaping mechanism is unclear. Therefore, using the modular components of ACEGEN we have conducted an ablation study to better understand the impact of the reward shaping, experience replay and sequence likelihood penalty that together comprise the components of the REINVENT algorithm. For this comparison, we used the default recommended REINVENT hyperparameters and measured performance on the MolOpt benchmark. With respect to maximum performance and efficiency, Table 4 shows that the primary performance improvement comes from the addition of experience replay to REINFORCE. While the addition of the reward shaping mechanism does improve over REINFORCE, the combination of experience replay and reward shaping does not improve over REINFORCE with experience replay. However, the addition of reward shaping that couples the agent policy to the prior policy does lead to an improved quality of chemistry as measured by the proportion of valid and unique molecules that also pass the chemistry filters, but at the compromise of performance. The results further emphasize the utility of adding regularisation in practical applications, as discussed in the previous section. In the original implementation, the authors of REINVENT also include a loss term to penalize highly likely sequences. Table 4 shows that this primarily increases exploration but to the detriment of all other metrics, including decreasing the proportion of molecules that pass chemistry filters, moreover, if used without reward shaping it leads to a complete collapse in learning.
@@ 3.4 Case Study: De-novo generation in the 5-HT2A
To reflect more realistic situations that may arise in practical drug discovery scenarios, we apply the different RL algorithms implemented on a more challenging objective previously proposed by Thomas et al. (2023). This objective is designing 5-HT2A receptor ligands selective over the highly related D2 receptor utilizing only structural information. This is an important and relevant challenge as clinically, this off-target profile leads to undesirable extrapyramidal side-effects in the treatment of psychosis Casey et al. (2022). To achieve this, this objective aims to minimize the docking score against 5-HT2A (PDB: 6A93), yet maximize the docking score against D2 (PDB: 6CM4), within a budget of 32,000 molecules. Both crystal structures are co-crystallized with Risperidone, reflecting the high degree of binding pocket similarity. In contrast to the proposed objective configuration, we use rDock Ruiz-Carmona et al. (2014) due to more permissive licensing.
@@
Table 5: Performance summary of algorithms on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 Gaulton et al. (2012). Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@
[…]
@@ 3.5 Case study: Scaffold constrained generation
ACEGEN can conduct constrained generation by leveraging PromptSMILES Thomas et al. (2024) which proposes iterative CLM prompting on SMILES in combination with RL to achieve fine-tune a pre-trained CLMs to the task of constrained generation. Here we demonstrate this application with two experiments: 1) scaffold decoration via known synthetic reactions as proposed with LibINVENT Fialková et al. (2021), and 2) scaffold decoration followed by scaffold-constrained docking to explore growth vectors inside a binding pocket.
In the first experiment, we compared the performance of RL algorithms on two tasks proposed with LibINVENT Fialková et al. (2021), both of which conduct scaffold decoration of a piperazine core with an amine linker moiety commonly present in D2 receptor ligands. As in the MolOpt benchmark, we apply a budget of 10,000 molecules. The first task is to optimize the predicted probability of D2 activity as evaluated by a QSAR model. The second task adds a selective reaction filter, rewarding piperazine growth via a Buchwald reaction and amine linker growth via an amide coupling reaction. Table 6 shows successful optimization of both objectives, with AHC performing best on the first task, and REINVENT (either with default or MolOpt parameters) performing best on the second task. However, all algorithms except A2C are able to solve the task with a yield above 90%, and fully satisfy the selective reaction objectives in 80 to 90% of molecules.
@@
In the second experiment, we used PromptSMILES in combination with AHC (based on the superior performance in the 5-HT2A case study) to conduct scaffold decoration in the catalytic site of BACE1, a key therapeutic target for the treatment of Alzheimer’s disease. Here we used the reference structure 4B05 co-crystallized with AZD3839 Jeppsson et al. (2012) which is a typical BACE1 inhibitor consisting of an amidine core interacting with two catalytic aspartates, and occupation of the P2’, P1 and P3 sub-pockets Oehlrich et al. (2014). The bicyclic, amidine-containing core was used as a prompt for PromptSMILES, allowing two growth vectors, one into the P2’ sub-pocket, and one into the P1 and P3 sub-pockets. Molecules were rewarded by minimizing the docking score using substructure constrained docking, as well as maintaining the number of heavy atoms and rotatable bonds in a sensible range (see Appendix F). Figure 2 shows the successful maximization of the reward over the generation of 10,000 molecules (Figure 2a,b), and corresponding minimization of the docking score (Figure 2c). The top 10 molecules identified contain improved docking scores in the range -23.9 to -15.8 compared to the reference molecule re-docked with a docking score of -8.44. Docked poses (shown in Figure 15) show the recovery of aromatic rings in the P1 and P2 sub-pockets. This case study highlights the successful constrained optimization of a complex objective.
@@ 4 Conclusion
In this study, we introduce ACEGEN, a novel toolkit that combines the best methodologies of reinforcement learning (RL) with the field of drug discovery. Leveraging RL building blocks from TorchRL, a general highly-tested decision-making library, ACEGEN provides modular, efficient and versatile solutions for drug discovery that we demonstrate with language-based models. We showcase ACEGEN’s capabilities across diverse areas, including method benchmarking, algorithmic exploration, and real-world drug discovery applications. Our experiments shed light on popular algorithms like REINVENT, offer practical insights into choosing the most suitable algorithm for specific contexts, and highlight the importance of dependable and comprehensive toolkits. Overall, ACEGEN addresses various needs in drug discovery and helps navigate the complex challenges associated with the field effectively.
@@
Appendix A Reinforcement Learning environment
@@
Appendix B Define scoring functions
@@
Appendix C Comparison against an existing implementation
We conducted a comparative analysis between our Reinvent implementation and MolOpt’s (Guo et al., 2022) implementation, which is accessible on their Github repository. To ensure a fair comparison, we replaced only the prior weights and utilized MolScore for computing scoring functions. Apart from these adjustments, we maintained the use of the original code. Both implementations use the same network architecture and hyperparameters.
@@ Appendix C Comparison against an existing implementation
Appendix D MolOpt Benchmark
The MolOpt benchmark was run using the MolScore implementation for convenience as it automatically calculates the same metrics as in the original implementation. In addition, MolScore computes the metrics after the removal of undesirable chemistry via the use of a basic chemistry filter (B-CF). This filter ensures molecules have a logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in Polykovskiy et al. (2020). If provided with a set of reference molecules, MolScore computes the metrics after the remove of undesirable chemistry via the use of a target chemistry filter (T-CF). This filter ensures molecules have a molecular weight and logP range within $\mu\pm 4\sigma$ of the reference distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). Lastly, MolScore also computes the metrics after filtering molecules using both basic and target chemistry filters (B&T-CF).
@@ Appendix D MolOpt Benchmark
Appendix E 5-HT2A case study
@@
A comparison of the protein interactions made by the top 10 compounds highlight the targeting of potentially selectivity inducing interactions. Figure 14 shows that many of the top 10 compounds proposed by ACEGEN-REINFORCE and ACEGEN-AHC form interactions with S131, T134, W151, and I152 in the second extended binding pocket which differ to those found in D2, and therefore, could potentially be a site to obtain selectivity. For example, occupied by the caged substructure shown in the selected examples in Figure 1 proposed by ACEGEN-AHC, occupying a sub-pocket that is not sterically possible in the D2 crystal structure used here. However, other D2 structures do show occupation of this after re-arrangement of extracellular loop 1 Fan et al. (2020). This reinforces the difficulty in identifying robust, validated scoring functions or oracles in drug discovery. For example, the use of Gypsum-DL Ropp et al. (2019) to prepare ligands for docking in this protocol leads to the generation of some implausible tautomers and unlikely protonation states and geometries (for example, cis-amides), meanwhile the docking parameters need further refinement to improve the quality of binding poses and cationic interactions with D1553x32 which are sometimes in an unlikely orientation. These can be swapped for commercial products like LigPrep in MolScore provided the licence is obtained. However, this challenging objective enables the differentiation of RL performance and behaviour in the optimization of a specified objective and sets a realistic and practical benchmark for drug discovery.
@@
Figure 9: Optimisation of a challenging 5-HT2A structure-based selectivity objective by different RL algorithms. The average score does not improve much over the course of training, but the average score of the top 100 compounds improves beyond known selective 5-HT2A ligands. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 Gaulton et al. (2012). Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@
Figure 14: Protein ligand interaction fingerprints of the top 10 compounds proposed by each RL algorithm. Examples are selected for visualization of their docking poses in Figure 1. Residue labels marked with an ”*” are different for the corresponding structurally aligned residue in D2.
Appendix F Scaffold constrained molecular generation case study
The same training dataset and pre-trained prior was used as for the MolOpt benchmark. In the first experiment, we used the objectives defined by LibINVENT Fialková et al. (2021) as re-implemented in MolScore. In the second experiment, we conducted a scaffold constrained docking experiment and took the product of three scores in the range [0-1] to define the final reward (see Table 15). The structure used was 4B05 Jeppsson et al. (2012). Note that during training, the inter-molecular rDock score was used as it was otherwise observed that RL would hack score optimization by improvement of intra-molecular rDock score only (data not shown). To discourage excessive exploitation, a diversity filter was used to penalize the re-generation of molecules via an ’occurance’ filter i.e., up to 5 duplicates resulted in no penalty, 5-10 duplicates resulted in an increasing penalty by multiplication of value scaling linearly from 1 to 0, and more than 10 duplicates resulted in a reward of 0. Both objectives and their configuration files to reproduce them can be found at https://github.com/MorganCThomas/MolScore/.

### v2

@@ 3.1 Benchmarking RL performance
To assess different RL algorithms, we have compared performance on the Practical Molecular Optimization (MolOpt) benchmark 10 as implemented in MolScore 39. This benchmark encompasses 23 distinct tasks associated with different objectives to be optimized within a budget of 10,000 molecules. All algorithms use the same GRU policy model architecture as in the original MolOpt benchmark, implemented in ACEGEN and pretrained on a curated subset of ChEMBL 32. We compared REINFORCE 19, REINVENT with the hyperparameters from the original paper 9, REINVENT-MolOpt with optimized hyperparameters for the MolOpt benchmark 10, and AHC with the hyperparameters from the original paper 6. Note that we utilize experience replay in all of the aforementioned algorithms Additionally, we also test A2C 20, PPO 21 which does not utilize any experience replay, and PPOD 22 which is an adaptation of PPO to accommodate experience replay. All algorithms are ACEGEN implementations, and the hyperparameter values used for all of them are provided in our repository. As a sanity check, we compared our implementation of REINVENT with the implementation used in the MolOpt paper. Our simplified re-implementation achieved better results with faster training. These results are shown in Appendix C C in the appendix.
@@
As proposed by the MolOpt benchmark authors, Table Table 1 shows algorithm performance on the AUC of the top 10 molecules as an indication of the sample efficiency of identifying 10 desirable molecules with respect to the objective, a representative figure that might be carried forward to later stages of drug design. This shows that PPOD is state-of-the-art with respect to sample efficiency, followed by REINVENT-MolOpt and then PPO. To achieve a higher level understanding of performance, Table 2 shows a selection of metrics aimed to measure maximum performance, efficiency, and exploration (a common trade-off with exploitation and hence efficiency in RL). These results show that REINVENT-MolOpt achieves maximum performance as measured by the average top 10 molecules, PPOD maximum efficiency as measured by AUC of the top 10 molecules, and AHC maximum exploration as measured by the number of unique compounds generated.
Table 2: Algorithm comparison on a selection of metrics analyzing the performance on the MolOpt benchmark. Valid is a sanity check that measures the proportion of valid molecules generated. Top-10 AUC is a measure of sample efficiency as shown in 1. Top-10 Avg is a measure of the absolute best 10 molecules achieved with the budget. Unique is a proxy for exploration which measures the proportion of unique molecules generated. Basic and target chemistry filters (B&T-CF) are the proportion of molecules after filtering out highly idiosyncratic molecules in general or with respect to the pretraining dataset. The B&T-CF Top-10 AUC and B&T-CF Top-10 Avg are recalculated on this subset while also enforcing that the 10 molecules identified offer diverse solutions (Div). Lastly, B&T-CF Diversity is a proxy measure of exploration via sphere exclusion diversity 44 on this subset. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. A perfect score for all metrics is 23.
@@ 3.2 Benchmarking RL performance for practical drug discovery
RL optimizes the policy for the cumulative future rewards as provided by either one scoring function or a sum of multiple scoring functions. However, in practical drug discovery, it is not straightforward to write the precise scoring functions that are needed and how to weigh them into a single scalar. For example, scoring functions can sometimes not produce the desired chemistry, or show clear exploitation loopholes 45, 46. In this setting, a sufficient element of regularization to a prior policy is useful, as the prior policy has learned a chemical space distribution based on a specified training dataset of choice which acts as a representation of desirable chemistry by examples. It is for this reason that methods like REINVENT have been designed heuristically to contain terms that try to enforce regularization instead of entirely using the reward signal.
To account for this, we make a modification to the average and AUC of the top 10 molecules to ensure they are diverse by an ECFP4 Tanimoto similarity of less than 0.35 to each other. Moreover, we use sphere exclusion diversity (SEDiv@1k) of de novo molecules 44 which measures the proportion of molecules needed to describe chemical space in a random sample of 1,000 molecules, as a proxy for exploration. This is more representative than the number of unique molecules that could all reside in a close area of chemical space. Lastly, we remove de novo molecules that do not pass a series of filters. This includes a basic chemistry filter (B-CF): logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 47. Moreover on the assumption that the training dataset describes desirable chemistry by examples, a target chemistry filter (T-CF): logP and molecular weight within $\mu\pm 4\sigma$ of the training dataset distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). The combination of the basic and target chemistry filter is denoted (B&T-CF).
@@ 3.3 Ablation study of the REINVENT algorithm
One of the most popular RL algorithms used for the fine-tuning of CLMs is REINVENT 4. From a theoretical RL perspective, this algorithm can be viewed a combination of the REINFORCE algorithm, experience replay, a reward shaping and a sequence likelihood penalty term for the loss function. However, the impact of each specific mechanism is unclear. Therefore, using the modular components of ACEGEN, we conducted an ablation study to better understand the impact of reward shaping, experience replay, and sequence likelihood penalty, which are key components of the REINVENT algorithm. We used default REINVENT hyperparameters and measured performance on the MolOpt benchmark for comparison. Regarding performance, Table 4 indicates that the primary improvement comes from adding experience replay to REINFORCE. While reward shaping does improve over REINFORCE, combining it with experience replay does not further improve performance. However, adding reward shaping that links the agent policy to the prior policy does improve the chemistry quality, measured by the proportion of valid and unique molecules passing the chemistry filters, albeit with a compromise in performance. These results highlight the importance of regularization in practical applications, as discussed in the previous section. In the original REINVENT implementation, the authors also include a loss term to penalize highly likely sequences. Table 4 shows that this primarily increases exploration but negatively affects all other metrics, including reducing the proportion of molecules passing chemistry filters. Furthermore, if used without reward shaping, it results in a complete collapse in learning.
@@ 3.4 Case Study: De-novo generation in the 5-HT2A
To reflect more realistic situations that may arise in practical drug discovery scenarios, we apply the different RL algorithms implemented on a more challenging objective previously proposed by 39. This objective is designing 5-HT2A receptor ligands selective over the highly related D2 receptor utilizing only structural information. This is an important and relevant challenge as clinically, this off-target profile leads to undesirable extrapyramidal side-effects in the treatment of psychosis 48. To achieve this, this objective aims to minimize the docking score against 5-HT2A (PDB: 6A93), yet maximize the docking score against D2 (PDB: 6CM4), within a budget of 32,000 molecules. Both crystal structures are co-crystallized with Risperidone, reflecting the high degree of binding pocket similarity. In contrast to the proposed objective configuration, we use rDock 49 due to more permissive licensing.
@@
Table 5: Performance summary of algorithms on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 32. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@
Figure 1: Selected examples from the top 10 molecules on the 5HT2A selective task and their docked pose in 5-HT2A (PDB: 6A93). The co-crystallised ligand Risperidone is included as the reference.
@@ 3.5 Case study: Scaffold constrained generation
ACEGEN can conduct constrained generation by leveraging PromptSMILES 43 which proposes iterative CLM prompting on SMILES in combination with RL to achieve fine-tune a pre-trained CLMs to the task of constrained generation. Here we demonstrate this application with two experiments: 1) scaffold decoration via known synthetic reactions as proposed with LibINVENT 18, and 2) scaffold decoration followed by scaffold-constrained docking to explore growth vectors inside a binding pocket.
In the first experiment, we compared the performance of RL algorithms on two tasks proposed with LibINVENT 18, both of which conduct scaffold decoration of a piperazine core with an amine linker moiety commonly present in D2 receptor ligands. As in the MolOpt benchmark, we apply a budget of 10,000 molecules. The first task is to optimize the predicted probability of D2 activity as evaluated by a QSAR model. The second task adds a selective reaction filter, rewarding piperazine growth via a Buchwald reaction and amine linker growth via an amide coupling reaction. Table 6 shows successful optimization of both objectives, with AHC performing best on the first task, and REINVENT (either with default or MolOpt parameters) performing best on the second task. However, all algorithms except A2C are able to solve the task with a yield above 90%, and fully satisfy the selective reaction objectives in 80 to 90% of molecules.
@@
In the second experiment, we used PromptSMILES in combination with AHC (based on the superior performance in the 5-HT2A case study) to conduct scaffold decoration in the catalytic site of BACE1, a key therapeutic target for the treatment of Alzheimer’s disease. Here we used the reference structure 4B05 co-crystallized with AZD3839 50 which is a typical BACE1 inhibitor consisting of an amidine core interacting with two catalytic aspartates, and occupation of the P2’, P1 and P3 sub-pockets 51. The bicyclic, amidine-containing core was used as a prompt for PromptSMILES, allowing two growth vectors, one into the P2’ sub-pocket, and one into the P1 and P3 sub-pockets. Molecules were rewarded by minimizing the docking score using substructure constrained docking, as well as maintaining the number of heavy atoms and rotatable bonds in a sensible range (see Appendix F). Figure 2 shows the successful maximization of the reward over the generation of 10,000 molecules (Figure 2a,b), and corresponding minimization of the docking score (Figure 2c). The top 10 molecules identified contain improved docking scores in the range -23.9 to -15.8 compared to the reference molecule re-docked with a docking score of -8.44. Docked poses (shown in Figure 15) show the recovery of aromatic rings in the P1 and P2 sub-pockets. This case study highlights the successful constrained optimization of a complex objective.
@@ 4 Conclusion
In this study, we introduce ACEGEN, a novel toolkit that combines the best methodologies of reinforcement learning (RL) within the field of drug discovery. Leveraging RL building blocks from TorchRL, a general highly-tested decision-making library, ACEGEN provides modular, efficient and versatile solutions for drug discovery that we demonstrate by implementing a suite of language-based solutions. We showcase ACEGEN’s capabilities across diverse areas, including method benchmarking, algorithmic exploration, and real-world drug discovery applications. Our experiments contribute to a better understanding of popular algorithms like REINVENT, offer practical insights into selecting the most suitable algorithm for specific contexts, and underscore the importance of dependable and comprehensive toolkits. Overall, ACEGEN addresses various needs in drug discovery and effectively navigates the complex challenges associated with the field.
@@
Appendix A A. Reinforcement Learning Environment
@@
Appendix B B. Define Scoring Functions
@@
Appendix C C. Comparison against an existing implementation
We conducted a comparative analysis between our Reinvent implementation and MolOpt’s 7 implementation, which is accessible on their Github repository. To ensure a fair comparison, we replaced only the prior weights and utilized MolScore for computing scoring functions. Apart from these adjustments, we maintained the use of the original code. Both implementations use the same network architecture and hyperparameters.
@@
Appendix D D. MolOpt Benchmark
The MolOpt benchmark was run using the MolScore implementation for convenience as it automatically calculates the same metrics as in the original implementation. In addition, MolScore computes the metrics after the removal of undesirable chemistry via the use of a basic chemistry filter (B-CF). This filter ensures molecules have a logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 47. If provided with a set of reference molecules, MolScore computes the metrics after the remove of undesirable chemistry via the use of a target chemistry filter (T-CF). This filter ensures molecules have a molecular weight and logP range within $\mu\pm 4\sigma$ of the reference distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). Lastly, MolScore also computes the metrics after filtering molecules using both basic and target chemistry filters (B&T-CF).
@@
Appendix E E. 5-HT2A case study
@@
A comparison of the protein interactions made by the top 10 compounds highlight the targeting of potentially selectivity inducing interactions. Figure 14 shows that many of the top 10 compounds proposed by ACEGEN-REINFORCE and ACEGEN-AHC form interactions with S131, T134, W151, and I152 in the second extended binding pocket which differ to those found in D2, and therefore, could potentially be a site to obtain selectivity. For example, occupied by the caged substructure shown in the selected examples in Figure 1 proposed by ACEGEN-AHC, occupying a sub-pocket that is not sterically possible in the D2 crystal structure used here. However, other D2 structures do show occupation of this after re-arrangement of extracellular loop 1 57. This reinforces the difficulty in identifying robust, validated scoring functions or oracles in drug discovery. For example, the use of Gypsum-DL 58 to prepare ligands for docking in this protocol leads to the generation of some implausible tautomers and unlikely protonation states and geometries (for example, cis-amides), meanwhile the docking parameters need further refinement to improve the quality of binding poses and cationic interactions with D1553x32 which are sometimes in an unlikely orientation. These can be swapped for commercial products like LigPrep in MolScore provided the licence is obtained. However, this challenging objective enables the differentiation of RL performance and behaviour in the optimization of a specified objective and sets a realistic and practical benchmark for drug discovery.
@@
Figure 9: Optimisation of a challenging 5-HT2A structure-based selectivity objective by different RL algorithms. The average score does not improve much over the course of training, but the average score of the top 100 compounds improves beyond known selective 5-HT2A ligands. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 32. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@
Figure 14: Protein ligand interaction fingerprints of the top 10 compounds proposed by each RL algorithm. Examples are selected for visualization of their docking poses in Figure 1. Residue labels marked with an "*" are different for the corresponding structurally aligned residue in D2.
Appendix F F. caffold constrained molecular generation case study
The same training dataset and pre-trained prior was used as for the MolOpt benchmark. In the first experiment, we used the objectives defined by LibINVENT 18 as re-implemented in MolScore. In the second experiment, we conducted a scaffold constrained docking experiment and took the product of three scores in the range [0-1] to define the final reward (see Table 15). The structure used was 4B05 50. Note that during training, the inter-molecular rDock score was used as it was otherwise observed that RL would hack score optimization by improvement of intra-molecular rDock score only (data not shown). To discourage excessive exploitation, a diversity filter was used to penalize the re-generation of molecules via an ’occurance’ filter i.e., up to 5 duplicates resulted in no penalty, 5-10 duplicates resulted in an increasing penalty by multiplication of value scaling linearly from 1 to 0, and more than 10 duplicates resulted in a reward of 0. Both objectives and their configuration files to reproduce them can be found at https://github.com/MorganCThomas/MolScore/.